OpenStudy (anonymous):

MY BRAIN ISN'T WORKING ANYMORE HELP!!!! Which of the following sets of numbers is not a Pythagorean triple? 3, 4, 5 10, 15, 20 5, 12, 13 7, 24, 25

OpenStudy (anonymous):

Which one of those sets of numbers makes the statement \(a^2+b^2=c^2\) false? Just plug them all in to find out which.

OpenStudy (anonymous):

3,4,5?

OpenStudy (anonymous):

Is the statement \(3^2+4^2=5^2\) true or false?

OpenStudy (anonymous):

false

OpenStudy (anonymous):

I am not sure witch it is ?

OpenStudy (anonymous):

How is \(3^2+4^2=5^2\) false? \(3\cdot3+4\cdot4=5\cdot5, 9+16=25\), how is that false?

OpenStudy (anonymous):

ooh! that makes sence! sorry

OpenStudy (anonymous):

So since that's true, that means that 3,4,5 is a pythagorean triple, which means it is not the answer. So now, find the one that is not a pythagorean triple.

OpenStudy (anonymous):

IDK

OpenStudy (anonymous):

I've already shown you exactly how to do everything that you need to do to solve this problem. Just plug in the rest of the numbers.

OpenStudy (anonymous):

thanks

OpenStudy (mertsj):

Did you find the answer?

OpenStudy (anonymous):

not realy!lol but I will figure it out some how! i am not that great at math

OpenStudy (mertsj):

Well, we know it is not the first one so let's try the second one. What is 10^2?

OpenStudy (anonymous):

thanks but i all ready turn it in but thanks for trying to help!lol

OpenStudy (mertsj):

yw